OpenStudy (anonymous):

will give medal and become fan - the base of a triangle is (8h + 14) centimeters. the height of the triangle is (2h - 9) centimeters. find the area of the triangle. A. 16h^2 - 44h - 126 B. 8h^2 - 44h - 63 C. 16h^2 - 100h - 126 D. 8h^2 - 22h - 63

OpenStudy (anonymous):

Area of a triangle formula = \[\frac{ 1 }{ 2 } \times b \times h\]

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

I got \[8x^2 - 22x - 63\]

OpenStudy (anonymous):

Well it's:\[8h^2 - 22h - 63\]

OpenStudy (anonymous):

@WeepingAngel123

OpenStudy (anonymous):

ok, so D?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

I couldn't really figure it out on paper, so I graphed it instead. It's 100% right.